# Training Neural Networks on Data Sources with Unknown Reliability

### v2

@@ A.8 Table 1 with multiple sources in batches
A.9 RRL + LAP with varied noise.
Table 5 shows the percentage difference in accuracy when using LAP over not using LAP for different noise levels and numbers of noisy sources.
In these experiments, data in CIFAR-10 is randomly split into $10$ sources, with the row of Table 5 defining the number of noisy sources, of which all have a noise rate as given by the column of Table 5. As before, the high capacity CNN with contrastive learning presented in Li et al. [16] is trained for $25$ epochs, with all other parameters kept as in the original work. Additionally, a version of this model (with the same parameters) is trained using LAP with $H=25$, $\delta=1.0$, and $\lambda=0.8$. These models are trained on batch sizes of $128$ using stochastic gradient descent with a learning rate of $0.02$, momentum of $0.9$, and weight decay $0.0005$. Data is randomly assigned to mini-batches, with each batch containing multiple sources.
When the whole dataset noise level is small (i.e: 2 sources with 25% noise), there are small performance increases when using LAP – likely because here, noisy sources have small impacts on the performance of models trained without LAP. Additionally, for lower noise levels, increasing the number of noisy sources reduced the performance improvement. This is likely because LAP is reducing noisy source training contributions early, when there is still information to learn. This can be remedied by increasing the leniency ($\lambda$), however since we were limited by compute, LAP parameters were fixed across experiments.
@@
Table 5: Different noise level and number of sources. Mean ± standard deviation (%) of the percentage difference in maximum test accuracy between RRL + LAP and RRL over $5$ repeats when training a model on CIFAR-10 for different noise levels and numbers of sources. Here a positive value represents an improvement in accuracy when using LAP. Here, $U$ corresponds to the number of unreliable sources out of 10 sources in total. Batch label flipping was used to introduce noise.
@@
A.10 CIFAR-10N with a smaller neural network
@@
In Figure 4, we show the results of the experiment presented in Figure 3(b) except when using the smaller network used to calculate the results in Table 1. The results in this experiment follow the same trend as in Figure 3(b), reassuring us of the applicability of LAP in a variety of settings, where lower capacity models are used.
@@
A.11 CIFAR-10 with a with large numbers of sources
@@
Table 6: Comparison of LAP with the baselines with varied numbers of sources. Mean ± standard deviation of maximum test accuracy (%) of $5$ repeats of the baselines and LAP on synthetic data with different noisy types and numbers of sources. The number of noisy sources ranges from $10$ to $50000$, corresponding to “LAP-$n$", where $n$ denotes the number of sources in the training data. For all experiments, $40\%$ of the sources were chosen as unreliable, with 100% noise rate. Only the LAP methods require re-training and so the results on the baselines are the same as in Table 4. Since there are $50000$ training data points, LAP-$50000$ corresponds to LAP acting over sources that have size $1$: i.e the standard noisy data setting.
@@
Model Types Noise Type Standard IDPA Co-teaching LAP-10 LAP-$1250$ LAP-$50000$ Original Data 69.46 ± 1.12 70.86 ± 0.32 67.73 ± 0.83 69.6 ± 1.08 69.57 ± 1.07 70.28 ± 0.5 Chunk Shuffle 65.92 ± 1.32 66.34 ± 0.99 63.55 ± 0.41 66.92 ± 0.56 66.08 ± 1.02 65.47 ± 0.98 Random Label 60.85 ± 1.67 57.93 ± 1.59 61.37 ± 1.01 66.5 ± 0.43 66.15 ± 1.16 66.49 ± 0.65 Batch Label Shuffle 62.23 ± 0.62 60.08 ± 0.13 63.9 ± 0.44 66.09 ± 0.65 66.71 ± 0.48 65.73 ± 1.13 Batch Label Flip 59.94 ± 2.48 61.02 ± 1.9 63.05 ± 2.43 64.48 ± 1.11 66.42 ± 1.22 65.41 ± 1.65 Added Noise 59.86 ± 1.38 61.19 ± 0.33 59.59 ± 1.37 63.58 ± 1.03 59.52 ± 1.06 60.55 ± 1.08 Replace With Noise 65.07 ± 0.63 65.85 ± 0.64 64.1 ± 1.13 64.89 ± 1.28 66.39 ± 1.66 65.09 ± 0.43
@@
To understand the effectiveness of our method as the number of sources grows, we extended the results in Appendix A.8 and Table 1 by significantly increasing the number of unique sources. In this experiment we see that LAP performs well across the different source sizes, demonstrating its robustness as the number of sources increases. In fact, LAP-$50000$ corresponds to an experiment where the size of each source is equal to $1$ – the standard noisy data setting. It is interesting to see that in this case LAP often performs as well or better than the baselines, suggesting its usefulness in a setting it was not originally designed for. All other experimental settings are kept the same from Table 1, with details in Appendix A.7.
These experiments take approximately $9$ hours to complete when using the compute described in Appendix A.1. This is because the results for the baselines are already calculated as part of the experiments in Appendix A.8.
A.12 Tiny-Imagenet results

### v4

@@ A.9 Experiments in detail
Results in Table 4. For this experiment, we train a ResNet 50 architecture using the baseline training method provided for Imagenet in Pytorch13. This trains the model for $90$ epochs with a batch size of $256$ using stochastic gradient descent with an initial learning rate of $0.1$, momentum of $0.9$, and weight decay of $0.0001$, as well as a learning rate scheduler that multiplies the learning rate by $0.1$ every $30$ epochs. We divide the Imagenet data into $10$ sources, of which $5$ are chosen to be noisy ($2$ containing input noise and $3$ containing label noise). The images are loaded as 8-bit integer arrays and interpolated to $(64,64)$ before the input noise is synthesised by randomly adding uniform integers from [-64, 64]. They are then transformed to 32-bit floats and normalised using the mean and standard deviation available on the same Pytorch training script. Label noise is added by randomly replacing labels in noisy sources. This dataset is then split into the ratio $9:1$ to produce a validation set for the analysis in Appendix A.18. A version of the ResNet 50 model (with the same parameters) is also trained using LAP with $H=25$, $\delta=1.0$ and $\lambda=0.8$. As before, data is randomly assigned to mini-batches, with each batch containing multiple sources. These experiments are repeated $5$ times for each setting.
A.10 Assessing the sensitivity of LAP to the hyperparameters
(a) Depression Strength, $\delta$.
(b) History Length, $H$.
(c) Leniency, $\lambda$.
Figure 6: Sensitivity to Hyperparameters. Here, we demonstrate the senstivitiy of the model performance based on the LAP hyperparamters introduced in Section 4 by varying their values whilst keeping the others fixed at $(\delta,H,\lambda)=(1.0,50,1.0)$.
In Figure 6 we present the change in cross entropy loss on the test set of the synthetic data used in Figure 4 and described in Section A.7. In this experiment the depression strength $\delta$, history length $H$, and leniency $\lambda$ were kept at $(\delta,H,\lambda)=(1.0,50,1.0)$ unless otherwise specified.
Depression strength. We observe that significantly increasing the depression strength $\delta$ can lead to a noticeable increase in test loss, particularly when $\delta\geq 2.0$ (Figure 6(a)). This is because large values of $\delta$ force training contributions from noisy data sources to be reduced early in training (as shown in Figure 4), whilst they might still be useful for learning a robust model. Suppressing them too quickly might prevent the model from learning important patterns in the noisy data which are useful for predicting on the test set, leading to worse performance. This highlights the need to choose a depression strength $\delta$ that strikes the right balance between filtering noise and learning from all available data. We find that a value of $\delta=1.0$ generally performs well.
History length. When studying the history length $H$, we see that the test loss remains relatively stable across the range of values (Figure 6(b)). There is some improvement for the history length values of $H=50$ and $H=100$, but there is little difference in performance between the use of large and small values of $H$. This indicates that whilst increasing $H$ allows for LAP to consider more contextual information when calculating the source weighting, it also makes our method slower to react to rapid changes in the training loss due to the significantly larger history being considered. We find that a $H$ value of $25$ to $50$ generally performs well.
Leniency. In Figure 6(c), we present the test loss as we vary the value of leniency $\lambda$. The loss values for this parameter remain relatively stable across its different values, suggesting that the model is fairly robust to variations in leniency. It may not be a critical parameter for tuning in this particular set up, since Figure 4 shows that for all leniency values tested the two noisiest sources were heavily weighted during training. In scenarios where noise levels are very low, it might be necessary to reduce the leniency to capture them. In contrast, when it is believed that patterns can be learnt from noisy data, it might be beneficial to use a larger leniency value, which intuitively lengthens the amount of time before noisy sources are weighted (Figure 4). In general, we find that a value of $\lambda=1.0$ performs well.
These experiments take approximately $5$ minutes to complete when using the compute described in Appendix A.1.
A.11 Table 1 with percentage difference in values
Table 6: Comparison of LAP with the baselines. Mean ± standard deviation of the percentage difference of the maximum test accuracy (%) of $5$ repeats of the baselines and LAP on synthetic data with different noisy types. For CIFAR-100 these numbers represent the top 5 accuracy. For CIFAR-10, CIFAR-100, and F-MNIST, the number of noisy sources are $4$, $2$, and $6$ out of $10$ respectively. Unreliable sources are each 100% noisy. All values in bold are within 1 standard deviation of the maximum score.
Model Types Noise Type Standard ARFL (2022) IDPA (2021) Co-teaching (2018) CDR (2021) LS (2022a) LAP (Ours) CIFAR-10 Original Data - -3.76% ± 2.40 2.66% ± 1.53 2.86% ± 0.45 -0.79% ± 2.03 0.84% ± 1.23 -0.38% ± 1.53 Chunk Shuffle - -5.12% ± 1.59 0.77% ± 0.49 3.68% ± 1.50 -2.89% ± 3.05 -2.79% ± 1.84 3.02% ± 2.22 Random Label - 3.10% ± 2.45 -17.90% ± 3.40 1.15% ± 1.43 -3.22% ± 3.44 0.15% ± 4.01 8.34% ± 3.93 Batch Label Shuffle - -2.31% ± 3.95 -0.80% ± 1.68 4.29% ± 2.08 -0.76% ± 2.18 -2.00% ± 2.74 7.59% ± 2.57 Batch Label Flip - -1.40% ± 3.67 3.77% ± 2.10 6.06% ± 1.05 -0.77% ± 3.30 1.90% ± 3.59 10.04% ± 0.85 Added Noise - -2.33% ± 1.51 1.02% ± 2.23 2.94% ± 0.91 -2.72% ± 2.79 -1.05% ± 2.28 3.43% ± 2.41 Replace With Noise - -7.92% ± 2.26 -1.12% ± 2.70 1.50% ± 0.42 -2.42% ± 1.50 0.26% ± 1.52 1.46% ± 1.63 CIFAR-100 Original Data - -20.42% ± 3.05 2.42% ± 0.77 0.41% ± 1.84 -0.02% ± 2.36 -0.67% ± 1.91 0.10% ± 1.91 Chunk Shuffle - -17.59% ± 3.31 -2.95% ± 1.83 1.06% ± 1.15 -0.29% ± 1.84 -1.03% ± 1.33 2.26% ± 1.72 Random Label - -16.20% ± 6.51 -15.13% ± 1.72 4.93% ± 1.96 -2.06% ± 1.89 -0.07% ± 0.90 18.40% ± 2.38 Batch Label Shuffle - -10.15% ± 7.25 -0.68% ± 3.85 7.03% ± 1.49 1.38% ± 1.39 -0.08% ± 2.65 6.96% ± 0.93 Batch Label Flip - -8.64% ± 6.29 4.41% ± 3.94 11.57% ± 2.15 -3.03% ± 2.58 -2.00% ± 3.63 12.84% ± 2.60 Added Noise - -10.60% ± 6.71 -1.27% ± 2.35 1.14% ± 1.34 -0.06% ± 2.33 0.15% ± 1.18 3.70% ± 2.24 Replace With Noise - -13.02% ± 7.46 -2.48% ± 2.26 1.12% ± 0.75 1.11% ± 3.49 -0.23% ± 2.76 1.46% ± 3.72 F-MNIST Original Data - -1.95% ± 0.83 0.07% ± 0.71 -5.44% ± 1.34 -1.44% ± 2.87 0.13% ± 0.67 -0.13% ± 0.31 Chunk Shuffle - 0.78% ± 2.35 0.47% ± 3.93 -3.06% ± 1.88 -0.81% ± 4.05 1.39% ± 2.17 6.61% ± 3.46 Random Label - 2.28% ± 11.45 0.75% ± 6.82 2.79% ± 15.20 0.30% ± 18.72 -0.85% ± 13.32 8.21% ± 10.42 Batch Label Shuffle - -4.29% ± 2.05 0.20% ± 0.57 -0.42% ± 0.29 0.28% ± 0.53 -0.02% ± 0.47 -0.24% ± 0.43 Batch Label Flip - -1.67% ± 4.39 2.61% ± 2.31 0.26% ± 2.64 1.33% ± 2.29 0.43% ± 2.28 -0.43% ± 2.86 Added Noise - -5.86% ± 3.02 0.77% ± 4.43 -2.71% ± 3.78 -2.09% ± 4.25 1.53% ± 3.13 1.14% ± 3.75 Replace With Noise - -7.07% ± 1.66 1.10% ± 1.08 -1.26% ± 1.20 -1.91% ± 1.20 0.22% ± 0.57 3.99% ± 1.05
Table 6 shows the values presented in Table 1 as a percentage difference from the standard training method. This more clearly demonstrates the size of the accuracy improvement from using LAP on CIFAR-10, CIFAR-100, and F-MNIST.
In particular, using LAP leads to substantial improvements over the baseline in more challenging scenarios, such as with random labelling noise or batch label flipping. For example, on CIFAR-100 with random label noise, LAP achieves a $21.02\%$ improvement in top-5 accuracy over standard training, significantly outperforming baselines such as IDPA. Even in cases where LAP is not the highest performing method, its performance remains comparable and often within a few percentage points of the best results.
This illustrates LAP’s robustness across varied datasets and noise types, showing that it consistently maintains or improves accuracy in high-noise conditions where other models struggle.
A.12 RRL + LAP with varied noise.
@@ A.11 Table 1 with percentage difference in values
Table 7: Different noise level and number of sources. Mean ± standard deviation (%) of the percentage difference in maximum test accuracy between RRL + LAP and RRL (2021) over $5$ repeats when training a model on CIFAR-10 for different noise levels and numbers of sources. Here a positive value represents an improvement in accuracy when using LAP. Here, $U$ corresponds to the number of unreliable sources out of 10 sources in total. Batch label flipping was used to introduce noise.
@@
Table 7 shows the percentage difference in accuracy when using LAP over not using LAP for different noise levels and number of noisy sources.
In these experiments, the data in CIFAR-10 is randomly divided into $10$ sources, the row of Table 7 defining the number of noisy sources, of which all have a noise rate as given by the column of Table 7. As before, the high capacity CNN with contrastive learning presented in Li et al. [17] is trained for $25$ epochs, with all other parameters kept as in the original work. In addition, a version of this model (with the same parameters) is trained using LAP with $H=25$, $\delta=1.0$, and $\lambda=0.8$. These models are trained on batch sizes of $128$ using stochastic gradient descent with a learning rate of $0.02$, momentum of $0.9$, and weight decay $0.0005$. Data is randomly assigned to mini-batches, with each batch containing multiple sources.
When the whole dataset noise level is small (i.e., 2 sources with 25% noise), there are small performance increases when using LAP – probably because here noisy sources have small impacts on the performance of models trained without LAP. Furthermore, for lower noise levels, increasing the number of noisy sources reduced the performance improvement. This is likely because LAP is reducing noisy source training contributions early, when there is still information to learn. This can be remedied by increasing the leniency ($\lambda$); however, since we were limited by the computation, the parameters of LAP were fixed across the experiments.
A.13 CIFAR-10N with a smaller neural network
@@
In Figure 7, we show the results of the experiment presented in Figure 5(b) except when using the low capacity CNN described in Appendix A.9. The results of this experiment follow the same trend as in Figure 5(b), which assures us of the applicability of LAP in a variety of settings, where models with lower capacity are used.
@@
A.14 CIFAR-10 with a with large numbers of sources
@@
Table 8: Comparison of LAP with the baselines with varied numbers of sources. Mean ± standard deviation of maximum test accuracy (%) of $5$ repeats of the baselines and LAP on synthetic data with different noisy types and numbers of sources. The number of noisy sources ranges from $10$ to $50000$, corresponding to “LAP-$n$”, where $n$ denotes the number of sources in the training data. For all experiments, $40\%$ of the sources were chosen as unreliable, with 100% noise rate. Since there are $50000$ training data points, LAP-$50000$ corresponds to LAP acting over sources that have size $1$: i.e the standard noisy data setting.
@@
Model Types Noise Type Standard IDPA (2021) Co-teaching (2018) CDR (2021) LS (2022a) LAP-10 LAP-$1250$ LAP-$50000$ Original Data 69.46 ± 1.12 70.86 ± 0.32 67.73 ± 0.83 70.88 ± 0.51 70.2 ± 0.66 69.6 ± 1.08 69.57 ± 1.07 70.28 ± 0.5 Chunk Shuffle 65.92 ± 1.32 66.34 ± 0.99 63.55 ± 0.41 66.18 ± 0.81 66.22 ± 0.84 66.92 ± 0.56 66.08 ± 1.02 65.47 ± 0.98 Random Label 60.85 ± 1.67 57.93 ± 1.59 61.37 ± 1.01 61.23 ± 1.16 61.1 ± 1.25 66.5 ± 0.43 66.15 ± 1.16 66.49 ± 0.65 Batch Label Shuffle 62.23 ± 0.62 60.08 ± 0.13 63.9 ± 0.44 64.63 ± 0.88 64.36 ± 1.02 66.09 ± 0.65 66.71 ± 0.48 65.73 ± 1.13 Batch Label Flip 59.94 ± 2.48 61.02 ± 1.9 63.05 ± 2.43 61.57 ± 1.06 61.84 ± 0.92 64.48 ± 1.11 66.42 ± 1.22 65.41 ± 1.65 Added Noise 59.86 ± 1.38 61.19 ± 0.33 59.59 ± 1.37 60.38 ± 0.84 60.05 ± 1.05 63.58 ± 1.03 59.52 ± 1.06 60.55 ± 1.08 Replace With Noise 65.07 ± 0.63 65.85 ± 0.64 64.1 ± 1.13 65.91 ± 1.03 65.7 ± 0.78 64.89 ± 1.28 66.39 ± 1.66 65.09 ± 0.43
To understand the effectiveness of our method as the number of sources grows, we extended the results in Table 1 by significantly increasing the number of unique sources. Here, we use the low capacity CNN presented in Appendix A.9, with all other settings kept the same as in Table 1, except that the number of epochs is reduced from $40$ to $25$ as the model architecture is smaller.
@@
In this experiment, we see that LAP performs well across the different source sizes, demonstrating its robustness as the number of sources increases. In fact, LAP-$50000$ corresponds to an experiment where the size of each source is equal to $1$ – the standard noisy datasetting. It is interesting to see that in this case LAP often performs as well or better than the baselines, suggesting its usefulness in a setting for which it was not originally designed.
A.15 Source distribution in GoEmotions dataset
The GoEmotions dataset [39] was chosen for its imbalanced source distributions in the real world. Here, the training set contains source sizes in the range of $1$ to $9320$ with a mean size of $1676$ and a standard deviation of $1477$, enabling us to study the robustness of LAP to imbalances in the source sizes and label distributions.
(a) The distribution of source sizes.
(b) Distribution of unique classes in sources.
Figure 8: Source and class distributions in GoEmotions. In 8(a) we present the number of sources with a given size, plotted on log-scale. This demonstrates the imbalance in the size of the sources in the dataset, with some containing thousands of data points, whilst others contain just a few hundred. In 8(b) we additionally explore the number of classes within each data source, showing that around half of all sources do not contain all classes, with two sources containing less than $5$ classes.
In Figure 8, we show the distribution of source sizes and the number of classes within each source.
Figure 8(a) demonstrates the imbalance in the size of each source, with some sources containing thousands of data points, whilst others contain hundreds or tens of data points. A dataset with this construction allows us to test the robustness of LAP to settings with uneven source sizes, a realistic situation in real-world data collection.
Furthermore, Figure 8(b) shows the distribution of the number of unique classes contained within each source. In particular, we observe that around half of all sources do not contain all the classes, with two sources containing fewer than $5$ classes. The imbalance in the class distribution across sources is more apparent when we list the largest three classes for some of the sources.
  • Source 1 with size $37$: class 28 = $40.54\%$, class 19 = $8.11\%$, class 26 = $5.41\%$
  • Source 2 with size $1137$: class 28 = $38.43\%$, class 19 = $10.82\%$, class 2 = $5.89\%$
  • Source 4 with size $2702$: class 5 = $17.84\%$, class 28 = $11.84\%$, class 11 = $8.62\%$
  • Source 8 with size $1253$: class 28 = $14.45\%$, class 23 = $12.05\%$, class 21 = $10.45\%$
  • Source 16 with size $3224$: class 5 = $13.43\%$, class 23 = $12.38\%$, class 1 = $10.92\%$
  • Source 32 with size $175$: class 28 = $50.86\%$, class 8 = $18.86\%$, class 2 = $6.86\%$
  • Source 64 with size $623$: class 28 = $18.46\%$, class 4 = $11.88\%$, class 5 = $6.58\%$
These values demonstrate that sources contain large variations in the number of data points and the distributions of classes they contain. By testing with this dataset we can verify LAP’s improved performance in settings in which the class distribution is uneven across sources, which may affect the source log-likelihood during training and strain the assumption we made that non-noisy sources contain similar data distributions.
A.16 Straining the assumptions of our method
To demonstrate the robustness of LAP, we now construct an experiment designed to strain the assumptions we made when introducing our method.
To work as intended, our proposed method assumes that all non-noisy sources contain similar data distributions. This allows us to say that the weighted likelihood ratio between a source under inspection and the other sources indicates whether to increase or decrease our reliability score (Equation 4). However, in real-world use cases, some sources might only produce a single class or a more challenging subset of classes. In such a case, we would like to understand whether our method can correctly identify the noisy sources without mistakenly labelling the more challenging data source as noisy.
To construct this setting, we combined data from both MNIST [48] and CIFAR-10 [32] (Appendix A.8). First, we evenly split the MNIST data into $99$ sources and applied random label noise to $95$ of them. This created a setting with many noisy sources, which we hypothesised would be a challenging setting to learn in. We then randomly chose $2$ classes from CIFAR-10 and assigned them to a single source, giving us $99$ MNIST sources (with $95$ of them noisy) and a single CIFAR-10 source (which remained non-noisy). We then kept the same non-noisy class labelling for MNIST from the original dataset (i.e.: in the non-noisy data, a handwritten digit 00 was assigned class 00, etc.) and mapped the two classes from CIFAR-10 to new classes: $10$ and $11$ – giving $12$ classes in total. Since CIFAR-10 is more challenging to classify than MNIST, this allowed us to construct a dataset in which almost all sources are easier to classify than the final source, which contains $2$ classes of more difficult data to separate. We also ran the same experiment with no random label noise applied to the MNIST sources.
To ensure that the CIFAR-10 and MNIST images are the same size and have the same number of filters, we applied a greyscale and resizing transformation to the CIFAR-10 data. We then use a simple CNN consisting of two convolutional layers with 32 and 64 filters, respectively, ReLU activation functions, max pooling layers (with kernel size $2$), and two lineaer layers mapping representations to sizes $9216$, $128$, and finally $13$. We use dropout, cross-entropy loss, and the Adam optimiser with learning rate $0.001$ and batches of size $128$, trained for $25$ epochs. The LAP parameters used were $(\delta,H,\lambda)=(1.0,50,4.0)$ for the original data experiment and $(0.1,50,1.0)$ for the random label experiment. These were chosen using $5$ runs on a validation set.
Table 9: Difficult data results. Mean ± standard deviation of maximum test accuracy (%) over $5$ repeats of standard training and LAP on a combination of MNIST and CIFAR-10 data with noisy labels. Here, $95$ out of $100$ sources are 100% noisy.
Model Types
Noise Type Standard LAP (Ours)
Original Data 98.38 ± 0.4 98.36 ± 0.39
Random Label 67.43 ± 2.83 96.29 ± 0.83
Table 9 shows the surprising results of this experiment. Since this data is made up of significant noise, it causes a large degradation in the accuracy of a standard training model when random labelling is applied to the data. However, LAP allows almost all of this loss in accuracy to remain, showing significantly higher performance on the test set than the standard training method, while allowing for comparative accuracy in the absence of noise. Here, we observe that even in a contrived experiment designed as a failure case for our proposed method, LAP still produces increased accuracy over this baseline.
Table 10: Difficult data results on the CIFAR-10 classes. Mean ± standard deviation of maximum test accuracy (%) on the CIFAR-10 classes over $5$ repeats of standard training and LAP trained on a combination of MNIST and CIFAR-10 data with noisy labels. This is the accuracy on just the CIFAR-10 test set, which is a subset of the test set corresponding to Table 9.
Model Types
Noise Type Standard LAP (Ours)
Original Data 0.97 ± 0.01 0.95 ± 0.06
Random Label 0.98 ± 0.01 0.97 ± 0.02
Furthermore, Table 10 shows that we do see a small reduction in performance on the more challenging source (containing the CIFAR-10 data). On the original data, in all five runs, LAP labelled the source containing the CIFAR-10 data as non-noisy as desired. However, when random label noise is applied to MNIST sources, the source containing CIFAR-10 data was incorrectly considered noisy in two of the five runs. In these two cases, the validation accuracy was lower than in the other three runs:
  • CIFAR-10 was incorrectly considered noisy: $23.50\%$ and $24.39\%$.
  • CIFAR-10 was correctly considered non-noisy: $25.46\%$, $25.19\%$, $25.53\%$.
Note that the validation accuracy is low because it contains the noisy labels as well as the clean labels. Therefore, with more runs and an ensemble of the higher validation accuracy runs, higher accuracy would likely be achieved on the test set and, in particular, on the CIFAR-10 source.
In the extreme case presented here, which violates the assumptions we made during Section 4, we find that our method is robust to challenging source class distributions and achieves markedly greater accuracy on the test set under noisy learning, whilst achieving comparative accuracy on the single challenging (but non-noisy) CIFAR-10 source.
(a) Training cross entropy loss.
(b) Validation top-5 accuracy.
(c) Test top-5 accuracy.
Figure 9: Performance on Imagenet. In 9(a) we show the cross-entropy loss on the training data at each epoch where the loss using LAP represents the tempered cross-entropy loss. In 9(b) we present the validation top-5 accuracy at each epoch. In 9(c) we present the test top-5 accuracy at each epoch to demonstrate the fitting to the noise that occurs when not using LAP during late training. Here, $5$ out of a total of $10$ sources are $100\%$ noisy, with $2$ containing input noise, and $3$ containing label noise. The two steps in performance occur at the points at which we scale the learning rate using a scheduler (epoch $30$ and $60$). The lines and error bands represent the mean and standard deviation over the $5$ repeats. The vertical black dashed line represents the epoch at which LAP achieved maximum top-5 accuracy on the test set.
A.17 Tiny-Imagenet results